In Hc.Verma concepts of phy vol-1 pg.72 ...
in the worked out examples problem no.1 how did they take the horizontal components and the vertical components .???

Can u explain this to me ???

I will b happy to approve plzzz..helpppp meee..


Aieee approaching givin me the nightmares pleaase helpppp

Always gonna be approved !!!
Thanks !@!

Grade:12

1 Answers

vikas askiitian expert
13 years ago

for this we have to consider a horizontal & vertical axis or x & y axis ...

tension can have 2 components , Tcos@ & Tsin@  ....

 

for left string tension has 2 components T2cosB along horizontal & T2sinB along verticle upward ...

for ryt string components are T1cos@ along horizontal & T1sin@ along verticle upward ...

 

now , at this stage we have to balance all the forces since we know that body is in equilibrium so no net force

exists in any direction ...

 

for upward direction , net force = T1sin@ + T2sinB

for downward dirctn , net force is weight of body = mg

for verticle equilibrium net force along verticle direction is 0 ..

Tsin@ + TsinB = mg          .......................1

 

for horizontal direction , T1cos@ = T2sinB        ...........2      (net force 0)

now solve 1 & 2 get the desired result

 

one more method can be used for all systems which are in equilibrium ,

lami's theorem =>

 the Lami's theorem is an equation relating the magnitudes of three coplanar, concurrent and non-collinear forces, which keeps an object in static equilibrium, with the angles directly opposite to the corresponding forces. According to the theorem,

\frac{A}{\sin \alpha}=\frac{B}{\sin \beta}=\frac{C}{\sin \gamma}
where A, B and C are the magnitudes of three coplanar, concurrent and non-collinear forces, which keeps the object in static equilibrium, and
α, β and γ are the angles directly opposite to the forces A, B and C respectively.
Lami's Theorem

Lami's theorem is applied in static analysis of mechanical and structural systems.